OpenStudy (anonymous):

the product of a number and - 7/10 is 1/2. what is the number -5/7 5/7 -7/20 7/20

OpenStudy (anonymous):

-7/20

OpenStudy (anonymous):

x*(-7/10)=1/2 times by -10/7 x=-5/7

OpenStudy (anonymous):

5/7

OpenStudy (anonymous):

-

OpenStudy (anonymous):

-7/20

OpenStudy (anonymous):

try pluging in what u got @Rolltideroll